OpenStudy (anonymous):

Write the expression (3 − 9i) + (4 + 5i) as a complex number in standard form.

OpenStudy (radar):

3 - 9i 4 +5i ------- 7 - 4i

OpenStudy (anonymous):

4-7i

OpenStudy (radar):

Is that theway you were told to do it?

OpenStudy (anonymous):

3-9i+4+5i=7-4i modulus=A=\[\sqrt{65}\] amrgument=t=arctan(-4/7)=pi-arctan(4/7) Z=A*exp(it)

OpenStudy (radar):

If you want it in polar form use vamgadu procedure.